solve for the values of x: (2x + 1) (x + 5) = (x + 11)( x + 3)
Answer by Boreal(15235)

Expand
2x^2+11x+5=x^2+14x+33
x^2-3x-28=0
(x-7)(x+4)=0
x=7, -4 ANSWER
15*12=18*10, true
-7*1=7*-1, also true